# Antisymmetric SU(2) adjoint generator as the universal origin of geometric phases in quantum two-level systems and classical polarization

### v2

@@ V Illustrative examples
V.1 Classical example: rotating wave plate on the Poincaré sphere
Consider a lossless ideal wave plate with retardance $\delta$ and fast axis defined by a unit vector $\mathbf{n}(\alpha)$ on the Poincaré sphere, where $\alpha$ denotes a rotation angle of the plate around the propagation direction. For definiteness, we take the fast axis to lie initially in the equatorial plane and to rotate around the $S_{3}$ axis, so that
$\mathbf{n}(\alpha)=\bigl(\cos 2\alpha,\;\sin 2\alpha,\;0\bigr)^{\mathsf{T}}.$ (34)
@@ V.1 Classical example: rotating wave plate on the Poincaré sphere
$U(\alpha)=\exp\!\left[-\frac{i}{2}\,\delta\,\mathbf{n}(\alpha)\cdot\bm{\sigma}\right],$ (35)
and the corresponding Poincaré-sphere rotation is given by Eq. (3) with $\mathbf{n}\to\mathbf{n}(\alpha)$. The antisymmetric generator is
$A(\alpha)=\sin\delta\,[\mathbf{n}(\alpha)]_{\times}=[\bm{\Omega}(\alpha)]_{\times},\qquad\bm{\Omega}(\alpha)=\sin\delta\,\mathbf{n}(\alpha).$ (36)
@@
As the plate rotates quasistatically from $\alpha=0$ to $\alpha=2\pi$ the Stokes vector $\mathbf{s}(\alpha)$ traces a closed curve $C$ on the Poincaré sphere. For a quarter-wave plate ($\delta=\pi/2$) this curve is a circle at a fixed latitude $\theta$ determined by the relative orientation between the input polarization and the eigenstates of the plate, while for a half-wave plate ($\delta=\pi$) it spans a larger portion of the sphere. In all cases, the instantaneous motion of $\mathbf{s}(\alpha)$ is governed by
$\frac{d\mathbf{s}(\alpha)}{d\alpha}=\bm{\Omega}(\alpha)\times\mathbf{s}(\alpha),$ (38)
@@
The explicit evaluation of $\Omega_{\mathrm{solid}}(C)$ depends on $\delta$ and on the initial state $\mathbf{s}_{\mathrm{in}}$, but its origin can be traced entirely to the antisymmetric generator $A(\alpha)$, since the symmetric part $S(\alpha)$ of the Mueller block does not enter the kinematic equation. For example, for a quarter-wave plate and a suitably chosen input state, the trajectory is a circle of colatitude $\theta$ on the sphere, and one obtains
@@
in agreement with the standard PB-phase result for such configurations. The present formalism shows that this phase can be read directly from $\bm{\Omega}(\alpha)$, which is determined by the antisymmetric part of the Mueller matrix of the rotating plate.
@@
V.2 Quantum example: spin-1/2 in a cyclic magnetic field
As a quantum counterpart, consider a spin-$1/2$ particle in a magnetic field of fixed magnitude $B$ whose direction precesses around the $z$ axis at a fixed polar angle $\theta$. The Hamiltonian is
@@ V.2 Quantum example: spin-1/2 in a cyclic magnetic field
$\mathbf{n}(t)$ $=\bigl(\sin\theta\cos\varphi(t),\;\sin\theta\sin\varphi(t),\;\cos\theta\bigr)^{\mathsf{T}}.$
where $\omega$ is the Larmor frequency and $\varphi(t)$ increases monotonically from 0 to $2\pi$ during a period $T$. The corresponding unitary evolution operator $U(t)$ is of the form (24), and its adjoint action on the Bloch vector is described by Eq. (25) with $\mathbf{n}\to\mathbf{n}(t)$.
@@
and $\delta(t)$ is related to the accumulated dynamical angle. The symmetric part $S_{U}(t)$ of the adjoint map does not contribute to $\gamma_{\mathrm{geom}}$.
@@
VI Operational aspects and possible extensions
@@
VI.1 Extraction from measured Mueller matrices
@@
is directly obtained and can be written as $A=[\bm{\Omega}]_{\times}$. The vector $\bm{\Omega}$ specifies the instantaneous angular velocity of the Stokes vector under the action of the retarder and, as shown in Sec. III, fully determines the geometric phase associated with any trajectory generated by the device.
@@
VI.2 Quantum process tomography and adjoint maps
In quantum information experiments, the evolution of a qubit is often characterized via quantum process to- mography [15, 17]. In the Bloch representation, a general quantum channel $\mathcal{E}$ acts as
@@ VI.2 Quantum process tomography and adjoint maps
coincides with the antisymmetric generator $A_{U}$ of Sec. IV. Therefore, once $R_{\mathcal{E}}$ has been reconstructed from experimental data, the geometric content of a unitary evolution is contained entirely in $A_{\mathcal{E}}$.
This observation suggests using the antisymmetric part of the reconstructed adjoint map as a diagnostic tool for geometric phases in noisy or imperfect implementations of qubit gates. In scenarios where $\mathcal{E}$ is close to a unitary channel, $A_{\mathcal{E}}$ provides an effective geometric generator that can be compared with the ideal design and used to detect deviations in the implemented geometric phase.
VI.3 Extensions beyond unitary two-level dynamics
@@
One direction is to consider more general completely-positive trace-preserving (CPTP) maps on qubits. In this case $R_{\mathcal{E}}$ need not be orthogonal, and the antisymmetric part $A_{\mathcal{E}}$ no longer generates a pure rotation. Nevertheless, $A_{\mathcal{E}}$ continues to encode the infinitesimal antisymmetric contribution to the mapping of Bloch vectors and may serve as a useful indicator of geometric features even in the presence of decoherence. A systematic analysis of geometric phases for nonunitary evolutions in terms of $A_{\mathcal{E}}$ lies beyond the scope of this work but appears as a natural extension.
Another direction is to explore higher-dimensional systems, such as spin-$j$ particles or multilevel atoms, whose pure-state manifolds can often be embedded in higher-dimensional analogues of the Bloch sphere. In these cases the relevant Lie algebras are $\mathfrak{su}(N)$ and their adjoint representations, which contain antisymmetric generators associated with rotations in generalized Bloch spaces. It is plausible that an appropriate generalization of the present construction identifies antisymmetric adjoint generators as carriers of geometric-phase information in those settings as well.
VII Conclusions
@@
These results establish a direct and operational bridge between classical polarization experiments, described by Mueller matrices, and quantum two-level dynamics, described by unitary adjoint maps on the Bloch sphere. In both cases, the complete geometric content of the evolution is encoded in the antisymmetric adjoint generator, which can be extracted from measured Mueller matrices or from quantum process tomography. Beyond its unifying role, this framework suggests concrete strategies for designing, controlling, and diagnosing geometric-phase effects in optical and quantum platforms, and it points naturally towards extensions to nonunitary channels and higher-dimensional systems where antisymmetric adjoint generators may play a similar geometric role.

### v3

@@ 5 Illustrative examples
5.1 Classical example: rotating wave plate on the Poincaré sphere
Consider a (lossless) ideal wave plate with retardance $\delta$ and fast axis defined by a unit vector $\mathbf{n}(\varphi)$ on the Poincaré sphere, where $\varphi$ is the azimuth of the fast-axis eigenpolarization in the laboratory frame. For definiteness, we take the fast axis to lie initially in the equatorial plane and to rotate around the $S_{3}$ axis, so that
$\mathbf{n}(\varphi)=\bigl(\cos 2\varphi,\;\sin 2\varphi,\;0\bigr)^{\mathsf{T}}.$ (36)
@@ 5.1 Classical example: rotating wave plate on the Poincaré sphere
$U(\varphi)=\exp\!\left[-\frac{i}{2}\,\delta\,\mathbf{n}(\varphi)\cdot\bm{\sigma}\right],$ (37)
and the corresponding Poincaré-sphere rotation is given by Eq. (3) with $\mathbf{n}\to\mathbf{n}(\varphi)$. The antisymmetric generator is
$A(\varphi)=\sin\delta\,[\mathbf{n}(\varphi)]_{\times}=[\bm{\Omega}(\varphi)]_{\times},\qquad\bm{\Omega}(\varphi)=\sin\delta\,\mathbf{n}(\varphi).$ (38)
@@
As the plate rotates quasistatically from $\varphi=0$ to $\varphi=2\pi$ the Stokes vector $\mathbf{s}(\varphi)$ traces a closed curve $C$ on the Poincaré sphere.
For a quarter-wave plate ($\delta=\pi/2$) and a fixed input state, the resulting trajectory is a closed, generally nonplanar loop on the Poincaré sphere. In the standard configuration of a fixed linear input polarization and a rotating linear retarder, this loop is the well-known three-dimensional figure-eight contour, given by the line of intersection of the Poincaré sphere with a right-circular cylinder whose axis is parallel to the $S_{3}$ direction [26]. For a half-wave plate ($\delta=\pi$) and the same fixed linear input, the trajectory remains confined to the equatorial plane (corresponding to linear polarization states), so that no solid angle is enclosed and the geometric phase vanishes.
In all cases, the instantaneous motion of $\mathbf{s}(\varphi)$ is governed by
$\frac{d\mathbf{s}(\varphi)}{d\varphi}=\bm{\Omega}(\varphi)\times\mathbf{s}(\varphi),$ (40)
@@
The explicit evaluation of $\Omega_{\mathrm{solid}}(C)$ depends on $\delta$ and on the initial state $\mathbf{s}_{\mathrm{in}}$, but its origin can be traced entirely to the antisymmetric generator $A(\varphi)$, since the symmetric part $S(\varphi)$ of the Mueller block does not enter the kinematic equation. For example, for a quarter-wave plate and a suitably chosen input state, the trajectory is a circle of colatitude $\theta$ on the sphere, and one obtains
@@
in agreement with the standard PB-phase result for such configurations. The present formalism shows that this phase can be read directly from $\bm{\Omega}(\varphi)$, which is determined by the antisymmetric part of the Mueller matrix of the rotating plate.
@@
5.2 Quantum example: spin-1/2 particle in a cyclic magnetic field
As a quantum counterpart, consider a spin-$1/2$ particle in a magnetic field of fixed magnitude $B$ whose direction precesses around the $z$ axis at a fixed polar angle $\theta$. Here the unit vector $\mathbf{n}(t)$ denotes the physical field direction in real space (laboratory coordinates), and the same $\mathbf{n}(t)$ enters the Bloch-sphere dynamics because the Hamiltonian generates rotations of the Bloch vector about the instantaneous field axis. The Hamiltonian is
@@ 5.2 Quantum example: spin-1/2 particle in a cyclic magnetic field
$\mathbf{n}(t)$ $=\bigl(\sin\theta\cos\phi(t),\;\sin\theta\sin\phi(t),\;\cos\theta\bigr)^{\mathsf{T}}.$
where $\omega$ is the Larmor frequency and $\phi(t)$ increases monotonically from 0 to $2\pi$ during a period $T$. The corresponding unitary evolution operator $U(t)$ is of the form (26), and its adjoint action on the Bloch vector is described by Eq. (27) with $\mathbf{n}\to\mathbf{n}(t)$.
@@
and in this setting the time-dependent rotation angle can be identified with the dynamical precession angle generated by the Hamiltonian, that is, with the time integral of the Larmor frequency for a field of fixed magnitude, so it plays the role of a rotation angle in Bloch space. The symmetric part $S_{U}(t)$ of the adjoint map does not contribute to $\gamma_{\mathrm{geom}}$.
@@
6 Operational aspects and possible extensions
@@
6.1 Extraction from measured Mueller matrices
@@
is directly obtained and can be written as $A=[\bm{\Omega}]_{\times}$. The vector $\bm{\Omega}$ specifies the instantaneous angular velocity of the Stokes vector under the action of the retarder and, as shown in Sec. 3, fully determines the geometric phase associated with any trajectory generated by the device. In realistic situations where the measured Mueller matrix deviates slightly from that of an ideal retarder due to weak diattenuation or depolarization, the antisymmetric part of the effective retarding block still provides a meaningful approximate geometric generator, capturing the dominant rotational contribution to the phase while the nonideal features mainly affect the symmetric and depolarizing components.
@@
6.2 Quantum process tomography and adjoint maps
In quantum information experiments, the evolution of a qubit is often characterized via quantum process tomography [20, 22]. In the Bloch representation, a general quantum channel $\mathcal{E}$ acts as
@@ 6.2 Quantum process tomography and adjoint maps
coincides with the antisymmetric generator $A_{U}$ of Sec. 4. Therefore, once $R_{\mathcal{E}}$ has been reconstructed from experimental data, the geometric content of a unitary evolution is contained entirely in $A_{\mathcal{E}}$.
This observation suggests using the antisymmetric part of the reconstructed adjoint map as a diagnostic tool for geometric phases in noisy or imperfect implementations of qubit gates. In scenarios where $\mathcal{E}$ is close to a unitary channel, $A_{\mathcal{E}}$ provides an effective geometric generator that can be compared with the ideal design and used to detect deviations in the implemented geometric phase. For example, in an experimental realization of a single-qubit geometric-phase gate in platforms such as superconducting circuits, trapped ions, or integrated photonics, one can reconstruct the process, extract the antisymmetric part of the adjoint map, and compare it directly with the ideal design to quantify geometric-phase errors even in the presence of moderate noise.
6.3 Extensions beyond unitary two-level dynamics
@@
One direction is to consider more general completely-positive trace-preserving (CPTP) maps on qubits. In this case $R_{\mathcal{E}}$ need not be orthogonal, and the antisymmetric part $A_{\mathcal{E}}$ no longer generates a pure rotation. Nevertheless, $A_{\mathcal{E}}$ continues to encode the infinitesimal antisymmetric contribution to the mapping of Bloch vectors and may serve as a useful indicator of geometric features even in the presence of decoherence. When the channel is close to unitary, this antisymmetric contribution can be interpreted as an effective geometric generator, whereas deviations from orthogonality in $R_{\mathcal{E}}$ and the accompanying translation vector reflect the nonunitary corrections induced by noise and decoherence. A systematic analysis of geometric phases for nonunitary evolutions in terms of $A_{\mathcal{E}}$ lies beyond the scope of this work but appears as a natural extension.
Another direction is to explore higher-dimensional systems, such as spin-$j$ particles or multilevel atoms, whose pure-state manifolds can often be embedded in higher-dimensional analogues of the Bloch sphere. In these cases the relevant Lie algebras are $\mathfrak{su}(N)$ and their adjoint representations, which contain antisymmetric generators associated with rotations in generalized Bloch spaces. It is plausible that appropriate generalizations of the present construction will identify antisymmetric adjoint generators as carriers of geometric-phase information in those settings as well.
7 Conclusions
@@
These results establish a direct and operational bridge between classical polarization experiments, described by Mueller matrices, and quantum two-level dynamics, described by unitary adjoint maps on the Bloch sphere. In both cases, the complete geometric-phase content of the evolution is encoded in the antisymmetric adjoint generator, which can be extracted from measured Mueller matrices or from quantum process tomography. The present framework suggests several directions for future work. On the classical side, the explicit identification of the antisymmetric Mueller generator as the carrier of geometric-phase information provides a natural tool to design and diagnose polarization-based geometric-phase elements, including space-variant retarders, geometric-phase lenses, metasurfaces, and birefringent fibers, directly from their measured Mueller matrices. For space-variant devices such as q-plates (i.e., birefringent elements whose optical axis varies across the transverse plane) and related subwavelength-grating elements, the analysis can be applied locally (pixel by pixel) to the position-dependent Mueller/Jones description, so that the spatially varying antisymmetric generator directly encodes the designed geometric-phase wavefront [24, 25]. On the quantum side, the same antisymmetric generator can be exploited to analyze the geometric content of single-qubit gates reconstructed via process tomography and to explore extensions to nonunitary channels and multilevel systems. In both regimes, the antisymmetric Mueller generator offers a compact language to interpret experimental data in terms of their underlying geometric structure.
Funding. No funding.
Disclosures. The author declares no conflicts of interest.
Data availability. No additional data were generated or analyzed in the course of this research.